OpenStudy (anonymous):

200 tickets were sold for the orchestra concert. Concert tickets were $15.00 for non-students and $11.00 for students. Total sales for the concert was $2680.00. How many tickets of each type were sold? A. n = 26, s = 174 B. n = 91, s = 109 C. n = 120, s = 80 D. n = 130, s = 70

hero (hero):

n + s = 200 15n + 11s = 2680 11n + 11s = 2200 15n + 11s = 2680 -11n - 11s = -2200 15n + 11s = 2680 4n = 480 n = 480/4 n = 120

OpenStudy (anonymous):

then do i plug it in for the s amount?

hero (hero):

200 - 120 = s

OpenStudy (anonymous):

thanks c:

hero (hero):

This is not difficult stuff

OpenStudy (anonymous):

It is! I skipped a math class and pretty much missed all of this soooo

hero (hero):

Well, you should have learned your lesson. Do not skip math classes.